Consider the experiment of tossing a coin three times (Prob. 1.1). Let X be the random variable giving the number of heads obtained. Assume that the tosses are independent and the probability of a head is p. (a) What is the range of X? (b) Find the probabilities P(X = 0), P(X = I), P(X = 2), and P(X = 3).


1
Expert's answer
2021-03-29T05:45:19-0400

A coin tossed three times. Total number of outcomes are "8."

Let "S" be sample space.

Then "S=" {"HHH,HHT,HTH,HTT,THH,THT,TTH,TTT" }

Let "X" be a random variable giving number of heads obtained.

Then "X" can take the values "0,1,2,3."

(a) The range of "X" is "=" {"0,1,2,3" }

(b) "P(X=0)=" probability of getting no head "=\frac{1}{8}"

"P(X=1)=" probability of getting one head "=\frac{3}{8}"

"P(X=2)=" probability of getting two heads "=\frac{3}{8}"

"P(X=3)=" probability of getting three heads "=\frac{1}{8}"